Need help in solving this equation by addition
2y + z = 4
x -3y + 2z =9
-y =3

Answer by Alan3354(69443)

2y + z = 4
x -3y + 2z =9
-y =3
--------------
y = -3
------
2y + z = 4
-6 + z = 4 Add 6 to both sides, that's addition
z = 10
---------
x +9 + 20 = 9
x = -20
-------